OpenStudy (anonymous):

2205 = 315a +210c 9 = a + c

OpenStudy (anonymous):

a=9-c 210c+315(9-c)=2205 -105c=-630 c=6 a=9-c a=3

OpenStudy (anonymous):

thanx